Circular No. 8418

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004ew AND 2004ex C. Jacques and E. Pimentel, Belo Horizonte, Brazil, report their dis-

covery of an apparent supernova (mag 17.3) on unfiltered CCD images taken on Oct. 8.18 and 10.15 UT. L. A. G. Monard, Pretoria, South Africa, reports his independent discovery of SN 2004ew at mag 17.5 on unfiltered CCD images taken on Oct. 9.918 and 10.758. Jacques provides the following position for the new object: α = 2^{h}05^{m}06^{s}.17, δ = −55^{o}06′31′′.6 (equinox 2000.0), which is 5′′.6 east and 12′′.1 north of the center of the galaxy ESO 153-G17. Jacques adds that nothing is visible at this position on a 1975 red Palomar Sky Survey image, Monard specifying that nothing is visible at this location on a red Digitized Sky Survey to limiting mag 20.5. Monard provides position end figures 06^{s}.16, 32′′.5. SN 2004ew also is not visible on earlier CCD images taken by the discoverers, with the following limiting magnitudes reported: 2003 Aug. 19.33, [18.5 (Jacques); 2004 Sept. 8.87, [18.5 (Monard); 12.21, [18.1 (Jacques).

Further to IAUC 7478, O. Trondal and M. Schwartz report the discovery of an apparent supernova (mag ∼ 17.7) on unfiltered Tenagra II 0.81-m telescope images taken on Oct. 11.34 and 13.33 UT. The new object was not detected in an earlier image taken on Sept. 24.27 (limiting mag ∼ 19.0). SN 2004ex is located at α = 0^{h}38^{m}10^{s}.19, δ = +2^{o}43′17′′.2 (equinox 2000.0), which is 33′′ west and 25′′.3 south of the nucleus of NGC 182.

SUPERNOVA 2004eu IN MCG +07-5-39 A. V. Filippenko, R. J. Foley, and B. J. Swift, University of Califor-

nia, Berkeley, report that inspection of CCD spectra (range 330–1000 nm), obtained on Oct. 13 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2004eu (IAUC 8417) is probably of type Ic, perhaps a week or two past maximum brightness. However, it exhibits some peculiarities, such as unusually weak Na I D absorption and unusually strong emission (Ca II?) at 400 nm.

M. Modjaz, R. Kirshner, P. Challis, Harvard-Smithsonian Center for Astrophysics; and T. Matheson, National Optical Astrononmy Observatory, report that a spectrum (range 350–740 nm) of SN 2004eu, obtained by P. Berlind on Oct. 8.37 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-Ic supernova. The spectrum is similar to spectra of SN 1994I (Filippenko et al. 1995, Ap.J. 450, L11), which were taken a few days after maximum. Follow-up spectra obtained by M. Calkin on Oct. 11.39 and 13.38 confirm the classification.

2004 October 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green